How many atoms of mercury are present in 5 cubic centimeters of liquid mercury? The desity of mercury is 13.55 g/cc
1
Expert's answer
2019-08-26T05:47:03-0400

"\bold{Given:}\\
Hg\\
V = 5 cm^3\\
\rho = 13.55 g/cm^3\\[1ex]

\bold{Find:}\\
N = \ ??\\[3ex]

N = \nu N_A = \dfrac{mN_A}{M} = \dfrac{\rho VN_A}{M} = \dfrac{13.55g/cm^3*5cm^3*6.02*10^{23}mol^{-1}}{200.59 g/mol} \approx 2.0333*10^{23}\\[1ex]
\bold{Answer:} \ 2.0333*10^{23} \ atoms"